OpenStudy (anonymous):

Twice the difference between a number and 4 is twenty. Find the number A.8 B.9 C.12 D.14

OpenStudy (anonymous):

The answer is D. 2x−8=20(Simplify both sides of the equation) 2x−8+8=20+8(Add 8 to both sides) 2x=28 (2x) 2 = 28 2 (Divide both sides by 2) x=14

OpenStudy (anonymous):

it is D